Place the following equation in the standard form of a quadratic equation.
6 + 4/x-1 = 5/x-3

Answer by edjones(8007)

6 + 4/(x-1) = 5/(x-3)
6(x-1)(x-3)+4(x-3)=5(x-1)
6(x^2-4x+3)+4x-12=5x-5
6x^2-24x+18+4x-12-5x+5=0
6x^2-25x+11=0
.
Ed